OpenStudy (anonymous):

A total of 428 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold . How many adult tickets were sold?

OpenStudy (anonymous):

EXPLAIN PLEASE

OpenStudy (anonymous):

A + S = 428 S = 3*A Solve.

OpenStudy (anonymous):

aaaaa!! idk

OpenStudy (anonymous):

Don't know how to solve: A + S = 428 S = 3A ?

OpenStudy (anonymous):

whats a?

OpenStudy (anonymous):

A is number of adult tickets. S is number of student tickets. The problem says total of them is 428. It also says 3 times more student tickets sold than adult tickets.

OpenStudy (anonymous):

soo... divide 428 by 3?

OpenStudy (anonymous):

No. If A +S = 428 and S = 3A, you substitute and get: A + 3A = 428

OpenStudy (anonymous):

107?

OpenStudy (anonymous):

Yes.

OpenStudy (cornzyblack):

first. Since the total tickets sold = 428 then, it means that the tickets were sold to both adults and Students So lets take A for adults and S for students so 428 = S + A But if number of students ticket was 3times that of adults then, S= 3A So, substitute 3A for S so 428 = 3A + A 428= 4A A = 107 Hence number of adult tickets = 107

OpenStudy (anonymous):

thanks guys!! muahhssss :*

OpenStudy (anonymous):

No medal?

OpenStudy (cornzyblack):

lol

OpenStudy (anonymous):

lol i didnt know i cang give medals !! lol